# SPECIAL RELATIVITY

### v11

@@ 5.4 A Simple Explanation of Light Aberration in Accelerated Frames
Above, we demonstrated that the metric tensor must be matched in both the accelerational and inertial segments of the trajectory. To achieve this, it is sufficient to relate the coordinates and time of the accelerated observer to those of the inertial observer using a Galilean boost. The electrodynamics equations in an accelerated frame can be obtained by applying a Galilean transformation (with velocity $-v$) to Maxwell’s equations. The electromagnetic fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame differ from the fields $\vec{E}$ and $\vec{B}$ in the inertial frame prior to the active Galilean boost. The transformation of the fields under the Galilean boost is given by $\vec{E}_{n}=\vec{E}-\vec{v}\times\vec{B}/c$, $\vec{B}_{n}=\vec{B}-\vec{v}\times\vec{E}/c$. This expression accounts only for first-order effects in $v/c$. Here, the velocity vector of the accelerated frame relative to the inertial frame is $\vec{v}=v\vec{e}_{x}$. If the electric field $\vec{E}$ is perpendicular to the $xz$-plane, the magnetic field in the inertial frame has only an $x$-component and we have immediately $\vec{E}_{n}=\vec{E}=\vec{e}_{y}E$, $\vec{B}_{n}=(\vec{e}_{z}\times\vec{e}_{y})E-(\vec{e}_{x}\times\vec{e}_{y})vE/c=-\vec{e}_{x}E-\vec{e}_{z}vE/c$. This results in a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$. Thus, the radiated beam propagates at an angle $-v/c$ with respect to the $z_{n}$-axis, leading to the phenomenon of aberration of light in the accelerated frame.
@@ 5.5 Clock Resynchronization in Accelerated Systems
Figure 16: Aberration of light in the accelerated frame of reference $S_{n}$. In this frame, the plane wavefront undergoes rotation following the re-synchronization of the accelerated clock. According to Maxwell’s electrodynamics, the transmitted light beam propagates at an angle of $-v/c$, resulting in the aberration of light.
The time coordinate $t^{\prime}_{n}$ in the $S_{n}$ frame, under Einstein synchronization, is obtained by introducing a time offset of the form given in Eq. 13, where the velocity $v$ is replaced with $-v$: specifically $-x_{n}v/c^{2}$. Applying this offset yields the first-order approximation $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$.
This time adjustment effectively rotates the plane of simultaneity by an angle $-v/c$. Consequently, after re-synchronization, the radiation wavefront undergoes rotation in the accelerated frame, as illustrated in Fig. 16.
The new time coordinate in the accelerated frame is interpreted in accordance with Maxwell’s equations, ensuring their applicability to light propagation. As a result, the transmitted light beam (passing through the aperture) propagates at an angle $-v/c$, demonstrating the phenomenon of the aberration of light. Both approaches yield the same outcome. The choice between these clock synchronization conventions is ultimately pragmatic. While changing the (four-dimensional) coordinate system does not introduce new physical phenomena, it provides a more consistent and coherent description of existing ones.
@@ 5.6 Composition of Motions in Non-Inertial Frames
It is always possible to choose such variables, in which the metric of an accelerated source will be diagonal. In the Lorentz coordinatization, light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source. If we look at the transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$. Substituting $dx_{n}/dt_{n}=-v$ into this equation confirms that $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$. The group velocity of the light beam remains $-v$, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$. This represents a correction to the classical aberration of light law within Lorentzian coordinates. Now, suppose an accelerated observer measures the aberration of light. Since time and distance in Lorentzian coordinates retain direct physical significance, the observer finds that the aberration angle satisfies the relation $\sin\theta_{a}=-v/c$.
@@ 5.7 Discussion
Where does the asymmetry originate? The electrodynamics equation must be integrated with initial conditions. According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation wavefront. In the inertial frame, the wavefront of the emitted light beam is perpendicular to the vertical $z$-axis. However, when considering Lorentz coordinatization, the wavefront orientation is affected in the accelerated frame $S_{n}$. According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration. Then, the radiated light beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
@@
Suppose (for simplicity) that there is a constant acceleration $g$ along the $x_{n}$-axis during the time $T=v/g$. The pseudo-gravitational acceleration is simply equal to the gradient of scalar potential: $g=-\partial\phi/\partial x_{n}$. A clock at a higher gravitational potential (aligned with the acceleration direction) runs faster. The pseudo-gravitational potential difference between two points along the $x_{n}$-axis is given by $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. Once the fixed stars move with constant velocity, the potential gradient in $S_{n}$ becomes zero. The accumulated time difference between two spatially separated clocks is $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$. This time shift effectively rotates the plane of simultaneity by an angle $-v/c$.
@@ 5.8 A Single Moving Emitter in an Accelerated Frame
We now consider the case where the emitter is at rest in the initial inertial frame $S$ (i.e., at rest relative to the fixed stars), while the observer, who is at rest in the accelerated frame $S_{n}$, performs the measurement of energy transport direction. It is important to emphasize that the aberration of light emitted by a single source also depends on the motion of the emitter within the accelerated frame $S_{n}$. Due to the fundamental asymmetry between inertial and accelerated frames, the theory of light aberration predicts a remarkable effect. Specifically, if the emitter remains at rest relative to the fixed stars while the observer transitions from rest to motion with respect to the fixed stars, the apparent angular position of the ”plane-wave” emitter, as seen in the accelerated frame, would jump by an angle of $-2v/c$. This aberration effect in the frame $S_{n}$ is illustrated in Fig. 17.
Let us describe the radiation emitted by a source moving with velocity $-v$ in an accelerated frame, as observed by an accelerated observer. The acceleration relative to the accelerated frame influences the propagation of light. Assuming an absolute time coordinatization, we apply a Galilean boost by substituting $x_{n}\to x_{n}+vt_{n}$ while keeping the time coordinate unchanged. Applying this transformation to the Langevin metric yields
@@
Another approach to solving the problem involves the Langevin metric diagonalization procedure. In this framework, the new coordinate system in the accelerated frame is interpreted such that an observer perceives a diagonal metric: $ds^{2}=c^{2}(dt^{\prime}_{n})^{2}-(dx_{n}^{\prime})^{2}-(dy_{n}^{\prime})^{2}-(dz_{n}^{\prime})^{2}$ for an emitter at rest in the accelerated frame, and $ds^{2}=c^{2}(1-v^{2}/c^{2})(dt^{\prime}_{n})^{2}-2vdx_{n}^{\prime}dt_{n}^{\prime}-(dx^{\prime}_{n})^{2}-(dy^{\prime}_{n})^{2}-(dz^{\prime}_{n})^{2}$ for a moving emitter. This reveals a symmetry (or reciprocity) in the metrics—and consequently in the electrodynamic equations—between the accelerated and inertial frames. This symmetry arises naturally from the pseudo-Euclidean geometry of space-time. The only difference between this metric and Eq. (11) is the sign of the velocity. After diagonalization, it becomes evident that any apparent asymmetry stems from the initial conditions. Notably, the plane radiation wavefront of a moving emitter (regardless of its velocity) undergoes a rotation by an angle of $-v/c$ in the accelerated frame as a result of the metric diagonalization.
To predict the outcome of aberration measurements, one can alternatively analyze the metric of the moving emitter directly within the absolute-time coordinate system, given by Eq. (18). According to electrodynamics, the group velocity $-2v$ emerges as a direct consequence of the cross term $-4vdxdt$.
@@ 5.10 Aberration of a Tilted Incoming Plane Wave
Now, let us examine the behavior of an incoming wave with a phase gradient along $x_{n}$. In other words, there is an angle $\alpha$ between the wavevector $\vec{k}$ and the $z_{n}$-axis. For small angles $\alpha\simeq v/c$, we can approximate $\alpha\simeq k_{x}/k_{z}$. Our goal is to express the parameters of the incoming wave in terms of the source offset $x_{o}$ relative to the optical axis. The wave tilt angle can be written as $\alpha\simeq x_{o}/z_{s}$, where $z_{s}$ is the distance from the source to the screen. Next, we analyze the aberration of light emitted by a stationary point source in an accelerated frame. The wavevector of the incoming plane wave is determined by the initial conditions. Assuming that the emitted wavefront is initially tilted by an angle $\alpha$, we obtain $k_{x}=\alpha k=\alpha\omega_{i}/c$, $k_{z}=\sqrt{\omega_{i}^{2}/c^{2}-k_{x}^{2}}$. Since acceleration modifies the field equations, we substitute the plane wave solution $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ into Eq. (17), yielding: $(1-v^{2}/c^{2})k_{x}^{2}-2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$. A screen with a hole acts as a diffraction grating, splitting the incident plane wave into multiple diffracted components. The wavenumber of the transmitted plane wave along the $x_{n}$-direction is $k_{x}=\alpha k+K_{\perp}$, where $K_{\perp}$ is the wavenumber associated with the sinusoidally modulated transmittance. The dispersion equation then simplifies to $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}\alpha kK_{\perp}$. This result indicates that the transmitted light beam propagates along $x_{n}$ with group velocity $d\omega/dk_{x}=-v+c\alpha$. Thus, the apparent angular position of the point source is related to the problem parameters by $\theta=-v/c+\alpha$  [^18] A similar effect arises if the screen moves with velocity $v$ along the $x$-axis in an inertial frame. In Chapter 4, we considered the special case of a normally incident plane wave, but this approach generalizes to tilted wavefronts as well. The deviation in energy transport for light passing through a hole in the moving screen leads to an aberration increment given by $\theta=v/c+\alpha$..
@@ 5.12 Two Metrics
On the other hand, the Langevin metric describes the electrodynamics of an accelerated emitter from the viewpoint of an accelerated observer. A key principle is that the metric tensor must remain continuous. This problem in special relativity can be effectively addressed using an absolute time coordinatization, which relies on Galilean boosts. To transform the inertial observer’s coordinates to the accelerated frame, we apply the inverse Galilean transformation: substituting $x\to x_{n}+vt,t\to t_{n}$ into the Minkowski metric yields the Langevin metric, Eq. (15). Notably, the Langevin metric Eq.15 of the accelerated emitter in the accelerated frame is the same as metric Eq.11 of the accelerated emitter in the inertial frame. The difference is only that the sign of velocity $v$. These two metrics are two sides of the same coin. Ultimately, we describe the electrodynamics of the same accelerated emitter—accelerated with respect to the fixed stars—from both an inertial and an accelerated observer’s viewpoint.
@@
In the problem under discussion, a puzzle remains. We introduce a comoving coordinate system in the initial inertial frame and analyze radiation from the moving source in terms of the new (comoving) coordinate labels. After performing a passive transformation, the source coordinates transform as $x^{\prime}=x-vt,t^{\prime}=t$. According to the equivalence of active and passive transformations within a single inertial frame, Maxwell’s equations remain valid in the comoving coordinate system. Consequently, the electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. This implies that the Minkowski metric, $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$, governs the electrodynamics of the accelerated light source in the comoving coordinate system. However, this raises a fundamental question: How can the Minkowski metric applies in the comoving coordinatization in the inertial frame while the Langevin metric applies in the accelerated frame? This apparent discrepancy requires further explanation.
@@
The difference between measurements made in the two reference frames is particularly intriguing. To analyze this, we first examine the relationship between passive and active transformations. It is important to note that a passive Galilean boost within a single inertial frame, given by $x^{\prime}=x-vt,t^{\prime}=t$, merely represents an alternative parametrization of the inertial observer’s measurements. Eq. (11) describes these measurements. In the case of active transformations within a single inertial frame, the motion of fixed stars relative to the inertial observer and their instruments remains unchanged. This establishes the equivalence between active and passive Galilean boosts. By contrast, the Langevin metric describes the measurements of an accelerated observer. Under the inverse Galilean boost, the motion of fixed stars relative to the space-time grid $(t_{n},x_{n},y_{n},z_{n})$ of the accelerated observer is altered.
On the other hand, the standard textbook treatment of Lorentz transformations implicitly assumes that the $x_{n}y_{n}z_{n}$-axes of the accelerated observer remain parallel to the $xyz$-axes of the inertial observer. In other words, it presumes that both observers share a common three-dimensional space. This assumption, often taken for granted, suggests that the simultaneous acceleration of a rigid reference frame has direct physical significance. However, this is a misconception. The conventional derivation of the aberration of light effect overlooks the conventionality of distant simultaneity. There is an inherent uncertainty in the relative position along the $z_{n}$-axis in the $x$-direction, amounting to $z_{n}v/c$, due to the ambiguity in the timing of acceleration. Consequently, we are dealing with space-like separated events. The orientation of the $z_{n}$-axis lacks an exact objective meaning because of the relativity of simultaneity. The claim that the $z_{n}$-axis orientation with respect to the $z$-axis of the inertial observer has objective meaning to within a certain accuracy can be visualized by the picture of $z_{n}$ axis in the vertical orientation with approximate angle uncertainty (blurring) given by $\Delta\theta=v/c$. This inclination angle between the inertial $z$ and accelerated $z_{n}$ reference directions has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this displacement could be ascertained. We will explore this issue further in Chapter 7.

### v12

@@ 5.4 A Simple Explanation of Light Aberration in Accelerated Frames
Above, we demonstrated that the metric tensor must be matched in both the accelerational and inertial segments of the trajectory. To achieve this, it is sufficient to relate the coordinates and time of the accelerated observer to those of the inertial observer using a Galilean boost. The electrodynamics equations in an accelerated frame can be obtained by applying a Galilean transformation (with velocity v-v) to Maxwell’s equations. The electromagnetic fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame differ from the fields $\vec{E}$ and $\vec{B}$ in the inertial frame prior to the active Galilean boost. The transformation of the fields under the Galilean boost is given by $\vec{E}_{n}=\vec{E}-\vec{v}\times\vec{B}/c$, $\vec{B}_{n}=\vec{B}-\vec{v}\times\vec{E}/c$. This expression accounts only for first-order effects in $v/c$. Here, the velocity vector of the accelerated frame relative to the inertial frame is $\vec{v}=v\vec{e}_{x}$. If the electric field $\vec{E}$ is perpendicular to the $xz$-plane, the magnetic field in the inertial frame has only an $x$-component and we have immediately $\vec{E}_{n}=\vec{E}=\vec{e}_{y}E$, $\vec{B}_{n}=(\vec{e}_{z}\times\vec{e}_{y})E-(\vec{e}_{x}\times\vec{e}_{y})vE/c=-\vec{e}_{x}E-\vec{e}_{z}vE/c$. This results in a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$. Thus, the radiated beam propagates at an angle v/c-v/c with respect to the $z_{n}$-axis, leading to the phenomenon of aberration of light in the accelerated frame.
@@ 5.5 Clock Resynchronization in Accelerated Systems
Figure 15: Aberration of light in the accelerated frame of reference $S_{n}$. In this frame, the plane wavefront undergoes rotation following the re-synchronization of the accelerated clock. According to Maxwell’s electrodynamics, the transmitted light beam propagates at an angle of v/c-v/c, resulting in the aberration of light.
The time coordinate $t^{\prime}_{n}$ in the $S_{n}$ frame, under Einstein synchronization, is obtained by introducing a time offset of the form given in Eq. 13, where the velocity $v$ is replaced with v-v: specifically $x_{n}v/c^{2}$. Applying this offset yields the first-order approximation $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$.
This time adjustment effectively rotates the plane of simultaneity by an angle v/c-v/c. Consequently, after re-synchronization, the radiation wavefront undergoes rotation in the accelerated frame, as illustrated in Fig. 15.
The new time coordinate in the accelerated frame is interpreted in accordance with Maxwell’s equations, ensuring their applicability to light propagation. As a result, the transmitted light beam (passing through the aperture) propagates at an angle v/c-v/c, demonstrating the phenomenon of the aberration of light. Both approaches yield the same outcome. The choice between these clock synchronization conventions is ultimately pragmatic. While changing the (four-dimensional) coordinate system does not introduce new physical phenomena, it provides a more consistent and coherent description of existing ones.
@@ 5.6 Composition of Motions in Non-Inertial Frames
It is always possible to choose such variables, in which the metric of an accelerated source will be diagonal. In the Lorentz coordinatization, light propagates with constant velocity $c$ independently of the direction of propagation, and of the velocity of its source. If we look at the transformation $t^{\prime}_{n}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x^{\prime}_{n}=x_{n}/\sqrt{1-v^{2}/c^{2}}$, we see that velocity along the $x_{n}$-axis is the same in the Lorentz coordinatization too: $dx_{n}^{\prime}/dt_{n}^{\prime}=(dx_{n}/dt_{n})/[1-v^{2}/c^{2}-v(dx_{n}/dt_{n})]$. Substituting $dx_{n}/dt_{n}=-v$ into this equation confirms that $dx_{n}^{\prime}/dt_{n}^{\prime}=-v$. The group velocity of the light beam remains v-v, but when we use the new velocity of light this becomes $\sin\theta_{a}=-v/c$. This represents a correction to the classical aberration of light law within Lorentzian coordinates. Now, suppose an accelerated observer measures the aberration of light. Since time and distance in Lorentzian coordinates retain direct physical significance, the observer finds that the aberration angle satisfies the relation $\sin\theta_{a}=-v/c$.
@@ 5.7 Discussion
Where does the asymmetry originate? The electrodynamics equation must be integrated with initial conditions. According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation wavefront. In the inertial frame, the wavefront of the emitted light beam is perpendicular to the vertical $z$-axis. However, when considering Lorentz coordinatization, the wavefront orientation is affected in the accelerated frame $S_{n}$. According to relativistic kinematics, the extra phase chirp $d\phi/dx^{\prime}_{n}=k_{x}=-v\omega_{i}/c^{2}$ is introduced as a consequence of this, the plane wavefront rotates after the acceleration. Then, the radiated light beam is propagated at the angle v/c-v/c with respect to the $z_{n}$-axis, yielding the phenomenon of the aberration of light in the accelerated frame $S_{n}$.
@@
Suppose (for simplicity) that there is a constant acceleration $g$ along the $x_{n}$-axis during the time $T=v/g$. The pseudo-gravitational acceleration is simply equal to the gradient of scalar potential: $g=-\partial\phi/\partial x_{n}$. A clock at a higher gravitational potential (aligned with the acceleration direction) runs faster. The pseudo-gravitational potential difference between two points along the $x_{n}$-axis is given by $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. Once the fixed stars move with constant velocity, the potential gradient in $S_{n}$ becomes zero. The accumulated time difference between two spatially separated clocks is $t_{n}(1)-t_{n}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$. This time shift effectively rotates the plane of simultaneity by an angle v/c-v/c.
@@ 5.8 A Single Moving Emitter in an Accelerated Frame
We now consider the case where the emitter is at rest in the initial inertial frame $S$ (i.e., at rest relative to the fixed stars), while the observer, who is at rest in the accelerated frame $S_{n}$, performs the measurement of energy transport direction. It is important to emphasize that the aberration of light emitted by a single source also depends on the motion of the emitter within the accelerated frame $S_{n}$. Due to the fundamental asymmetry between inertial and accelerated frames, the theory of light aberration predicts a remarkable effect. Specifically, if the emitter remains at rest relative to the fixed stars while the observer transitions from rest to motion with respect to the fixed stars, the apparent angular position of the ”plane-wave” emitter, as seen in the accelerated frame, would jump by an angle of 2v/c-2v/c. This aberration effect in the frame $S_{n}$ is illustrated in Fig. 16.
Let us describe the radiation emitted by a source moving with velocity v-v in an accelerated frame, as observed by an accelerated observer. The acceleration relative to the accelerated frame influences the propagation of light. Assuming an absolute time coordinatization, we apply a Galilean boost by substituting $x_{n}\to x_{n}+vt_{n}$ while keeping the time coordinate unchanged. Applying this transformation to the Langevin metric yields
@@
Another approach to solving the problem involves the Langevin metric diagonalization procedure. In this framework, the new coordinate system in the accelerated frame is interpreted such that an observer perceives a diagonal metric: $ds^{2}=c^{2}(dt^{\prime}_{n})^{2}-(dx_{n}^{\prime})^{2}-(dy_{n}^{\prime})^{2}-(dz_{n}^{\prime})^{2}$ for an emitter at rest in the accelerated frame, and $ds^{2}=c^{2}(1-v^{2}/c^{2})(dt^{\prime}_{n})^{2}-2vdx_{n}^{\prime}dt_{n}^{\prime}-(dx^{\prime}_{n})^{2}-(dy^{\prime}_{n})^{2}-(dz^{\prime}_{n})^{2}$ for a moving emitter. This reveals a symmetry (or reciprocity) in the metrics—and consequently in the electrodynamic equations—between the accelerated and inertial frames. This symmetry arises naturally from the pseudo-Euclidean geometry of space-time. The only difference between this metric and Eq. (11) is the sign of the velocity. After diagonalization, it becomes evident that any apparent asymmetry stems from the initial conditions. Notably, the plane radiation wavefront of a moving emitter (regardless of its velocity) undergoes a rotation by an angle of v/c-v/c in the accelerated frame as a result of the metric diagonalization. This implies that the radiated light beam has a group velocity along $x$-direction given by $d\omega/dk_{x}=-2v$.
To predict the outcome of aberration measurements, one can alternatively analyze the metric of the moving emitter directly within the absolute-time coordinate system, given by Eq. (18). According to electrodynamics, the group velocity 2v-2v emerges as a direct consequence of the cross term 4vdxdt-4vdxdt.
@@ 5.10 Aberration of a Tilted Incoming Plane Wave
Now, let us examine the behavior of an incoming wave with a phase gradient along $x_{n}$. In other words, there is an angle $\alpha$ between the wavevector $\vec{k}$ and the $z_{n}$-axis. For small angles $\alpha\simeq v/c$, we can approximate $\alpha\simeq k_{x}/k_{z}$. Our goal is to express the parameters of the incoming wave in terms of the source offset $x_{o}$ relative to the optical axis. The wave tilt angle can be written as $\alpha\simeq x_{o}/z_{s}$, where $z_{s}$ is the distance from the source to the screen. Next, we analyze the aberration of light emitted by a stationary point source in an accelerated frame. The wavevector of the incoming plane wave is determined by the initial conditions. Assuming that the emitted wavefront is initially tilted by an angle $\alpha$, we obtain $k_{x}=\alpha k=\alpha\omega_{i}/c$, $k_{z}=\sqrt{\omega_{i}^{2}/c^{2}-k_{x}^{2}}$. Since acceleration modifies the field equations, we substitute the plane wave solution $\exp(i\vec{k}_{n}\cdot\vec{r}_{n}-i\omega_{n}t_{n})$ into Eq. (17), yielding: $(1-v^{2}/c^{2})(k_{n})_{x}^{2}-2v(k_{n})_{x}\omega_{n}/c^{2}+(k_{n})_{z}^{2}-\omega_{n}^{2}/c^{2}=0$. A screen with a hole acts as a diffraction grating, splitting the incident plane wave into multiple diffracted components. The wavenumber of the transmitted plane wave along the $x_{n}$-direction is $(k_{n})_{x}=\alpha k+K_{\perp}$, where $K_{\perp}$ is the wavenumber associated with the sinusoidally modulated transmittance. The dispersion equation then simplifies to $\omega_{i}\Delta\omega=-\omega_{i}K_{\perp}v+c^{2}\alpha kK_{\perp}$. This result indicates that the transmitted light beam propagates along $x_{n}$ with group velocity $d\omega/dk_{x}=-v+c\alpha$. Thus, the apparent angular position of the point source is related to the problem parameters by $\theta=-v/c+\alpha$  [^18] A similar effect arises if the screen moves with velocity $v$ along the $x$-axis in an inertial frame. In Chapter 4, we considered the special case of a normally incident plane wave, but this approach generalizes to tilted wavefronts as well. The deviation in energy transport for light passing through a hole in the moving screen leads to an aberration increment given by $\theta=v/c+\alpha$..
@@ 5.11 Non-Reciprocity in the Aberration of Light Theory
To further explore the aberration of light in an accelerated frame, it is useful to consider a slightly modified source—a ”plane wave” emitter. We adopt the model described in Section 4.2 because it is both relatively simple and sufficiently general to serve as a prototype for understanding aberration phenomena. While no one has performed all the ”thought experiments” exactly as described here, the outcomes can be confidently predicted based on the laws of special relativity, which are themselves grounded in experimental evidence.
@@ 5.12 Two Metrics
On the other hand, the Langevin metric describes the electrodynamics of an accelerated emitter from the viewpoint of an accelerated observer measurements. A key principle is that the metric tensor must remain continuous. This problem in special relativity can be effectively addressed using an absolute time coordinatization, which relies on Galilean boosts. To transform the inertial observer’s coordinates to the accelerated frame, we apply the inverse Galilean transformation: substituting $x\to x_{n}+vt,t\to t_{n}$ into the Minkowski metric yields the Langevin metric, Eq. (15). Notably, the Langevin metric Eq.15 of the accelerated emitter in the accelerated frame is the same as metric Eq.11 of the accelerated emitter in the inertial frame. The difference is only that the sign of velocity $v$. These two metrics are two sides of the same coin.
@@
In the problem under discussion, a puzzle remains. We introduce a comoving coordinate system in the initial inertial frame and analyze radiation from the moving source in terms of the new (comoving) coordinate labels. After performing a passive transformation, the source coordinates transform as $x^{\prime}=x-vt,t^{\prime}=t$. According to the equivalence of active and passive transformations within a single inertial frame, Maxwell’s equations remain valid in the comoving coordinate system. Consequently, the electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. This implies that the Minkowski metric, $ds^{2}=c^{2}(dt^{\prime})^{2}-(dx^{\prime})^{2}-(dy^{\prime})^{2}-(dz^{\prime})^{2}$, governs the electrodynamics of the accelerated light source in the comoving coordinate system.
However, a conceptual tension emerges: Why does the Minkowski metric apply in the comoving description (within an inertial frame), while the Langevin metric governs the physics in the truly accelerated frame? This apparent inconsistency requires careful examination.
The key lies in recognizing that inertial and accelerated observers do not share the same three-dimensional space in Minkowski spacetime. Although this might seem paradoxical if both observers were assumed to share a common spatial geometry, special relativity resolves the issue: observers following different worldlines inhabit different instantaneous 3-spaces due to the relativity of simultaneity.
To analyze this, we begin by examining the relationship between passive and active transformations. A passive Galilean boost within a single inertial frame, given by $x^{\prime}=x-vt,t^{\prime}=t$, simply re-expresses the results of measurements made by the inertial observer using a new set of coordinates.
@@
In the case of active transformations within the same inertial frame, the motion of fixed stars relative to the observer and their instruments remains unchanged. This establishes the equivalence between active and passive Galilean boosts in inertial frames.
In contrast, the Langevin metric describes measurements made by an accelerated observer. Under an inverse Galilean boost, the motion of fixed stars relative to the accelerated observer’s space-time grid $(t_{n},x_{n},y_{n},z_{n})$ is no longer preserved. This highlights the fundamental difference in how motion is perceived in non-inertial frames.
When observers adopt absolute-time coordinatization, relativistic kinematics becomes remarkably simple. In this approach, all asymmetries between the initial inertial and accelerated frames are absorbed into their respective frame metrics.
Alternatively, consider observers using Lorentz coordinatization. Here, the frame metrics become diagonal, and the complexities of special relativity manifest as counterintuitive kinematic effects. Notably, in this coordinatization, the inertial and accelerated observers experience different three-dimensional spatial geometries.
On the other hand, the standard textbook treatment of Lorentz transformations implicitly assumes that the $x_{n}y_{n}z_{n}$-axes of the accelerated observer remain parallel to the xyzxyz-axes of the inertial observer. In other words, it presumes that both observers share a common three-dimensional space. This assumption, often taken for granted, suggests that the simultaneous acceleration of a rigid reference frame has direct physical significance. However, this is a misconception.
For example, conventional derivation of the aberration of light effect overlooks the conventionality of distant simultaneity. There exists an inherent ambiguity in the spatial position along the $z_{n}$-axis (in the $x$-direction) due to uncertainties in the timing of acceleration. This uncertainty can be quantified as $z_{n}v/c$, revealing that the associated events are space-like separated.
Consequently, the orientation of the $z_{n}$-axis lacks an exact objective meaning, due to the relativity of simultaneity. The idea that the direction of the $z_{n}$-axis relative to the inertial $z$-axis is physically well-defined is, at best, an approximation. This relationship can be visualized as a vertical $z_{n}$ axis with an angular uncertainty (or ”blurring”) given by: $\Delta\theta=v/c$. This inclination angle has no precisely defined physical meaning, since—given the finite speed of light—there is no experimental method to determine such a spatial displacement.
Let us illustrate these ideas with the example of a plane-wave emitter moving along the $x$-axis with velocity $v$ in an initial inertial frame. Assume that clocks are synchronized according to the Einstein synchronization procedure, using light signals from a stationary dipole source. The electrodynamics of the moving source in this inertial frame is described by the metric in Eq. 11. As previously shown, the radiated beam exhibits aberration: due to the motion of the source, the direction of energy propagation is altered.
However, when we introduce a comoving coordinate system via a passive Galilean boost, this aberration disappears. In this comoving system, the light beam propagates along the $z$-axis. Crucially, because the boost is performed within a single inertial frame, the comoving coordinate axes $(x^{\prime},y^{\prime},z^{\prime})$ remain parallel to the original inertial axes $(x,y,z)$. That is, both systems inhabit the same three-dimensional space.
Let us now examine the relationship between the coordinate system of the inertial frame and that of the comoving frame. To detect aberration in the inertial frame, one must first establish a coordinate system, which requires a physical, operational procedure. In this sense, the spatial grid $(x,y,z)$ of the inertial frame represents a physical reality. By contrast, the comoving coordinate grid $(x^{\prime},y^{\prime},z^{\prime})$ introduced through the passive Galilean boost is a purely mathematical construct—there is no need to consider how it would be physically realized.
It is important to emphasize that a passive Galilean boost within a single inertial frame merely provides an alternative parametrization of the same observations made by an inertial observer using real, physical reference axes.
Now consider the perspective of an accelerated observer. As before, to detect any aberration effect within the accelerated frame, a coordinate system must be established. This requires the observer to define a practical, operational method for assigning spatial coordinates $(x_{n},y_{n}z_{n})$.
The accelerated observer finds that the angular displacement of the beam is negative, given by $\theta_{a}=-v/c$. This result leads to a key insight: from the viewpoint of the inertial observer, the $z_{n}$-axis of the accelerated observer is not parallel to the inertial $z$-axis. In other words, the inertial and accelerated observers inhabit different three-spaces within Minkowski spacetime.
This subtle but important distinction reflects the breakdown of simultaneity in relativistic physics and anticipates deeper discussions that will follow in Chapter 7.